# RAT [Uncaptioned image]: Bridging RNN Efficiency and Attention Accuracy in Language Modeling

### v1

@@ 4 Related work
RNN
Recurrent neural networks (RNNs) have a long history of modeling sequential data [28, 29, 30], but suffer from training instabilities and lack parallelizability, especially when handling long sequences. Various approaches have been proposed to improve their trainability, such as gating mechanisms [10, 11], unitary designs [31, 32], and partially parallelizable variants [33, 34]. Moreover, RNNs face a memory bottleneck as they compress entire sequences into fixed-size hidden-state vectors. Some earlier work addressed this issue by allowing the current step to access earlier time steps directly [35, 36, 37]. For example, Koutnik et al. [36] introduces multiple timescales by updating hidden states only at predefined clock rates, while Ke et al. [37] proposes sparse, learnable connections to past tokens, with weights computed by a small MLP. More recently, state space models [38, 12, 39, 13, 40] have emerged, expanding hidden states from vectors to matrices through state expansion and reduction techniques. They also replace standard matrix multiplications with structured operations (e.g. diagonal matrices) to further enable parallel computation [14]. In contrast to these approaches, and inspired by the success of attention mechanisms, we propose integrating attention directly over the RNN history to allow explicit retrieval of past information. This design remains simple and naturally interpolates between RNN and attention behaviors. By applying inter-chunk attention to retrieve information, we avoid state expansion, which may introduce training instability and typically requires careful normalization or initialization [12, 5].
Attention
While attention mechanisms have proven highly effective, they suffer from slow computation due to their full access to all previous tokens, especially in long sequences. Recently, there has been a growing interest in linear attention [4, 6, 41, 13, 5, 42, 26], which explores the dual formulation of linear RNN and linear attention. These approaches rely on strict linearity constraints and introduce state expansion and per-head gating techniques. In contrast to this duality perspective, we adopt an intermediate view between per-dimension-gated RNNs and softmax-based attention, without imposing hard linearity constraints on either side. The use of linear recurrence in our model is primarily motivated by training efficiency, but it is not strictly required at the design level. Other works have explored attention within local contexts, often motivated by the diagonal structures observed in attention maps, and have introduced recurrence or sparse attention mechanisms to provide access to earlier context [1, 2, 43, 44, 3, 45]. For example, Ma et al. [43] applies attention in chunks after gating the query and key vectors, and Behrouz et al. [3], inspired by cognitive science, focuses attention on the short-range context. In our work, we take a different approach: we use attention to model long-range dependencies, leveraging their strength in directly accessing distant tokens, while applying RNNs to short-range contexts, where they are typically less prone to memory degradation. This can also be seen as a form of attention with resolution, loosely mirroring how humans process information: short-term inputs are integrated into coherent representations, while long-term events are stored as memory anchors that can be selectively retrieved when relevant. We also find that RAT and the attention of sliding windows [1] are structurally complementary. The interleaving of the two leads to further performance improvements while being efficient.
@@ 5 Conclusion
Limitations. Due to resource constraints and the exploratory nature of this work, our experiments are limited to models with up to 1.3B parameters. We have not yet scaled RAT to larger language models such as 7B or 14B to confirm our results hold at that scale. Additionally, we did not adopt supervised fine-tuning techniques commonly used in current industry practices, which typically involve substantial data resources and many engineering details. Instead, we followed a more classical approach based on train-test splits to evaluate performance, as our focus lies primarily on architectural design during pretraining. Lastly, RAT may still face length generalization issues from positional encoding, similar to attention. But as shown in Subsec. C.2, this can be significantly reduce by shorter inter-chunk attention spans, and we also explore the use of NoPE.
@@ A.2 Experiments
In our preliminary studies, we use the PG19 dataset [46], a long-form English book corpus with inherently long contexts. For the 1.3B model experiments, we adopt the FineWeb-Edu dataset [20], using its 100B-token randomly sampled version downloaded from the HuggingFace repository. To match the pretraining context length, we concatenate documents using a separator token. Note that web samples are usually very short, compared to the book dataset. For downstream evaluation, we consider a suite of classical commonsense reasoning benchmarks from the Eleuther AI evaluation harness [22], including PIQA [47], ARC-C [48], and HellaSwag [49]. In the LLaMA2 tokenizer, we observe that the inputs of these tasks typically contain fewer than 300 tokens. For the LongBench evaluation, in LABEL:tab:longbench we provide a mapping of task numbering to names and annotate their input lengths for each task, offering a rough indication of task difficulty with respect to input length. For SFT-based tasks, we have elaborated sequence lengths in the main text. As LongBench and SFT tasks often involve very long inputs, we apply truncation in the middle to preserve information at both the beginning and the end.
Table 5: We provide a mapping from task numbers to task names. NQA (NarrativeQA), MQA (MultiFieldQA-en), HQA (HotpotQA), WQA (2WikiMultihopQA), MSQ (Musique), GR (GovReport), MN (MultiNews), TQA (TriviaQA), and RBP (RepoBench-P). We also report the average input length and the 95th percentile input length of each task, measured in tokens using the LLaMA2 tokenizer.
Task Single-Document QA Multi-Document QA Summarization Few-shot learning Code Completion
1-1 1-2 1-3 2-1 2-2 2-3 3-1 3-2 3-3 4-1 4-2 4-3 6-1 6-2
Name NQA Qasper MQA HQA WQA MSQ GR QMSum MN TREC TQA SAMSum LCC RBP
Avg. 36037 5780 8115 15329 8483 18555 12280 15980 3156 7792 14098 11168 4307 14818
95th pctl. 77966 10164 14994 19755 16939 20066 25721 29069 7135 12532 25006 19823 10401 31937
@@
We start with a 200M-parameter model in our preliminary study, with a model dimension of 1024, 12 transformer layers, and head dimension of 64. The rotary position embedding (RoPE) base is set to 10,000. We use the GPT2 tokenizer. Following Mohtashami and Jaggi [45], we repeat the PG19 training split five times to reach a total of 15B training tokens. The learning rate is scheduled using cosine annealing, starting at $6.0\times 10^{-4}$ and decaying to $1.0\times 10^{-6}$, with a warm-up ratio of 10%. We use the AdamW optimizer with a weight decay of 0.1 and $\beta=(0.9,0.98)$. Gradient clipping is applied with a threshold of 1.0, and the global batch size is set to 1M tokens. We explore training with three different context lengths: 8K, 16K, and 32K. Training these models on 4 H100 GPUs takes approximately 5 to 14 hours. In particular, the attention model requires up to 14 hours when the sequence length is $T=32768$, whereas the RAT(L=16) model completes training in about 7 hours under the same setting.
@@
The 1.3B-parameter model uses a model dimension of 2048, 24 transformer layers, and a head dimension of 128, equipped with RMSNorm [50] and without bias. The RoPE base is also set to 10,000. The model parameters are initialized using a Gaussian distribution with a standard deviation of 0.02. We adopt the LLaMA2 tokenizer in the following studies. For pretraining, we use a cosine-annealed learning rate schedule starting at $8.0\times 10^{-4}$ and decaying to $1.0\times 10^{-6}$, with 5% warmup. The global batch size is set to 2M tokens. We primarily train with a context window of 4K tokens and include additional experiments with a 16K context window, for which the RoPE base is increased to 500,000. Each model is trained on 16 H100 GPUs, requiring approximately 2 to 3 days to complete.
For LongBench evaluation, we follow the default prompts with greedy decoding for all tasks except summarization. For summarization, we apply a repetition penalty of 1.2 to address the common issue in pretrained-only models of generating repetitive outputs under instructional prompts. For SFT tasks, we train the models on the official training splits with an answer-only loss and evaluate them on the corresponding test sets. Although we explored different hyperparameters during the early experimentation, we observed that the relative trends remained largely stable. Thus, we fix the learning rate and batch size to $(1.0\times 10^{-5},128)$ for large datasets, and $(1.0\times 10^{-5},32)$ for the smaller QMSum [24] task, following common practice for 1B-scale models. The weight decay is set as 0.01, and all other hyperparameters follow the pretraining setup. We sweep over the number of epochs and report the best result for each dataset and architecture. In practice, QA tasks typically converge in 1–2 epochs, while summarization benefits from slightly longer training.

### v3

@@ 5 Related work
State space and linear attention models
Recurrent neural networks (RNNs) have long been used for sequential modeling [elman1990finding, siegelmann1992computational, bengio1994learning]. Recent state space models [s4, s5, mamba, griffin, lru] extend recurrence by expanding hidden states from vectors to matrices for higher accuracy, removing non-linearity and using structured operations (e.g., diagonal matrices) to enable parallel computation [s5]. Per-head gating connects these models [mamba2] to linear attention [linear_attention, retnet, rwkv, lru, deltanet, gateddeltanet] and shows the duality between linear recurrence and linear attention. Unlike this duality perspective, we take an intermediate view between per-dimension-gated RNNs and softmax-based attention, without enforcing strict linearity. Linear recurrence in RAT is intended primarily for training efficiency, rather than imposed as a design constraint.
Despite state expansion, these models inherit the fixed-size representation limitation of classic RNNs, leading to degraded memory on long sequences. A concurrent work [log_linear_attention] proposes addressing this by growing the memory slots logarithmically. We similarly target the fixed-size memory problem, but use softmax-based attention to access distant tokens. Hybrid strategies [samba, hymda] combining state space models and attention are also popular. While RAT is not a hybrid with the same computation graph for all tokens and all layers, it is orthogonal to such designs and provides more flexibility, such as using different chunk sizes for different layers and heads, which we leave for future work.
Softmax-based attention
Attention mechanisms suffer from slow computation due to their full-token access. Earlier works have explored local attention with chunking, combined with recurrence or sparse mechanisms to access earlier context [longformer, gau, mega, megalodon, landmarkattention]. For example, mega gates query/key vectors before local attention, while gau adds the outputs of local attention to cross-chunk linear attention. In our work, we instead leverage attention’s strength in directly accessing distant tokens with lightweight recurrence for local contexts, and hierarchically organize their outputs. As shown in Subsec. C.1, RAT utilizes FLOPs more effectively than local attention. We also regard RAT and local attention as structurally complementary. Although we experimented with interleaving them in this work, we leave the combination with more advanced local attention variants in future research.
RAT can be viewed as an attention mechanism with resolution, akin to how humans process information: short-term inputs are integrated into coherent representations, while long-term events are stored as memory anchors for selective retrieval. Compared to dilated attention [longformer], our simple recurrence enables global perception, and the chunk-level attention has high training parallelism and reduced cache memory (see Subsec. C.1 for more). Finally, a recent work [gatedslotattention] also aggregates the key and value vectors and applies softmax-based attention over them. However, it compresses the full-sequence vectors into a 2D memory matrix, unlike our chunk-based design, which enables flexible memory slots.
@@ 6 Conclusion
Limitations. Due to resource constraints and the exploratory nature of this work, our experiments are limited to models with up to 1.3B parameters. We have not yet scaled RAT to larger language models such as 7B or 14B to confirm our results hold at that scale. Additionally, we did not adopt supervised fine-tuning techniques commonly used in current industry practices, which typically involve substantial data resources and many engineering details. Instead, we followed a more classical approach based on train-test splits to evaluate performance, as our focus lies primarily on architectural design during pretraining. Lastly, RAT may still face length generalization issues from positional encoding, similar to attention. But as shown in Subsec. C.2, this can be significantly reduced by shorter inter-chunk attention spans, and we also explore the use of NoPE.
Acknowledgments and Disclosure of Funding
We sincerely thank the RCP and the SCITAS team at EPFL for GPU support. We also thank the Swiss AI Initiative and the Swiss National Supercomputing Centre (CSCS) for supporting this work through grants under project IDs a06 and a10. We thank Skander Moalla, Anja Surina, and Lars Quaedvlieg for helpful discussion on SFT tasks. We extend our appreciation to Karin Getaz for administrative support.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We ensure that the abstract and introduction clearly summarize the paper’s core contributions: the RAT architecture, its intra- and inter-chunk mechanisms, the efficiency, and empirical evaluations on both pretraining and downstream tasks. All major claims are supported by experiments and analyses throughout the paper.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We discuss limitations in Sec. 6
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: The paper does not include theoretical results.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We describe our experimental setup and implementation details in Sec. 4 and Appendix A.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We open-source our code with link put in Appendix. The dataset we use is publicly available.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We describe our experimental setup in Sec. 4 and give details in Appendix A.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: We do not report error bars or statistical significance due to computational cost and limited variance observed across repeated runs in preliminary testing. We follow the common practice in similar-scale LLM experiments.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Computing resources are discussed in Sec. 4 and Appendix A.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The research conducted in the paper conforms with the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: Societal impacts are discussed in Appendix D.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: We do not include safeguards in the current version, as model and data release is not the focus of this work, and the model is only 1B scale. Should we release models in the future, we will follow community standards to mitigate misuse risk.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: All the assets used have been cited.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We provide documented code.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: LLMs were not used in any component of the core research methodology. Any usage was limited to writing assistance and did not affect the technical contributions of the work.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
@@ A.2 Experiments
In our preliminary studies, we use the PG19 dataset [pg19], a long-form English book corpus with inherently long contexts. For the 1.3B model experiments, we adopt the FineWeb-Edu dataset [fineweb_edu], using its 100B-token randomly sampled version downloaded from the HuggingFace repository. To match the pretraining context length, we concatenate documents using a separator token. Note that web samples are usually very short, compared to the book dataset. For downstream evaluation, we consider a suite of classical commonsense reasoning benchmarks from the Eleuther AI evaluation harness [eval-harness], including PIQA [piqa], ARC-C [arc], and HellaSwag [hellaswag]. In the LLaMA2 tokenizer, we observe that the inputs of these tasks typically contain fewer than 300 tokens. For the LongBench evaluation, in LABEL:tab:longbench we provide input length for each task, offering a rough indication of task difficulty with respect to input length. For SFT-based tasks, we have elaborated sequence lengths in the main text. As LongBench and SFT tasks often involve very long inputs, we apply truncation in the middle to preserve information at both the beginning and the end.
Table 7: We report the average input length and the 95th percentile input length of each task, measured in tokens using the LLaMA2 tokenizer. NQA (NarrativeQA), MQA (MultiFieldQA-en), HQA (HotpotQA), WQA (2WikiMultihopQA), MSQ (Musique), GR (GovReport), MN (MultiNews), TQA (TriviaQA), and RBP (RepoBench-P).
Task Single-Document QA Multi-Document QA Summarization Code Completion
Name NQA Qasper MQA HQA WQA MSQ GR QMSum MN LCC RBP
Avg. 36037 5780 8115 15329 8483 18555 12280 15980 3156 4307 14818
95th pctl. 77966 10164 14994 19755 16939 20066 25721 29069 7135 10401 31937
@@
We start with a 200M-parameter model in our preliminary study, with a model dimension of 1024, 12 transformer layers, and head dimension of 64. The rotary position embedding (RoPE) base is set to 10,000. We use the GPT2 tokenizer. Following landmarkattention, we repeat the PG19 training split five times to reach a total of 15B training tokens. The learning rate is scheduled using cosine annealing, starting at $6.0\times 10^{-4}$ and decaying to $1.0\times 10^{-6}$, with a warm-up ratio of 10%. We use the AdamW optimizer with a weight decay of 0.1 and $\beta=(0.9,0.98)$. Gradient clipping is applied with a threshold of 1.0, and the global batch size is set to 1M tokens. We explore training with three different context lengths: 8K, 16K, and 32K. Training these models on 4 H100 GPUs takes approximately 5 to 14 hours. In particular, the attention model requires up to 14 hours when the sequence length is $T=32768$, whereas the RAT(L=16) model completes training in about 7 hours under the same setting.
@@
The 1.3B-parameter model uses a model dimension of 2048, 24 transformer layers, and a head dimension of 128, equipped with RMSNorm [rmsnorm] and without bias. The RoPE base is also set to 10,000. The model parameters are initialized using a Gaussian distribution with a standard deviation of 0.02. We adopt the LLaMA2 tokenizer in the following studies. For pretraining, we use a cosine-annealed learning rate schedule starting at $8.0\times 10^{-4}$ and decaying to $1.0\times 10^{-6}$, with 5% warmup. The global batch size is set to 2M tokens, and the context window is set to 4096. Each model is trained on 16 H100 GPUs, requiring approximately 2 to 3 days to complete.
For LongBench evaluation, we follow the default prompts with greedy decoding for all tasks except summarization. For summarization, we apply a repetition penalty of 1.2 to address the common issue in pretrained-only models of generating repetitive outputs under instructional prompts. For SFT tasks, we train the models on the official training splits with an answer-only loss and evaluate them on the corresponding test sets. Although we explored different hyperparameters during the early experimentation, we observed that the relative trends remained largely stable. Thus, we fix the learning rate and batch size to $(1.0\times 10^{-5},128)$ for large datasets, and $(1.0\times 10^{-5},32)$ for the smaller QMSum [qmsum] task, following common practice for 1B-scale models. The weight decay is set as 0.01, and all other hyperparameters follow the pretraining setup. We sweep over the number of epochs and report the best result for each dataset and architecture. In practice, QA tasks typically converge in 1–2 epochs, while summarization benefits from slightly longer training.